Need help in solving this question

  • Allocate resource between two time periods t=1,2
  • The fixed supply of the resource is 10 barrels 
  • Demand is same in both periods and given by Pt = 15 - qt
  • Marginal cost of extracting the resource is the same in both periods and is given by MCt=$3
  • The discount rate is 7%

1. What is the dynamically efficient allocation of this resource in period 1? Please enter this number rounded to the hundredths place in the 1st box below

1
Expert's answer
2021-03-18T16:29:18-0400
"Solution"

"P-MC(Q1)=\frac{[P-MC(Q2)]}{1+r}\\"


given : "P=15-Qt\\"

"MC=\$3"

"Q1+Q2=10"

"15-3+3.5Q1=\frac{[15-(3+3.5Q2)}{1+0.07}"


"13+3.5Q1=\frac{[13-3.5Q2]}{1.07}"


"Q2=10-Q1"


"13-3.5Q1=\frac{13}{1.07}-\frac{3.5(10-Q1)}{1.07}"


"13-\frac{13}{1.07}+\frac{35}{1.07}=\frac{3.5Q1+3.5Q2}{1.07}"


"33.5607=\frac{7Q1}{1.07}"


"Q1=\frac{33.5607}{6.5420}"


"Q1=5.130"


for period 1 is "5.130"